find the line that passes through the point (1,2) and is parallel to the line x+2y=3
Answer by Alan3354(69443)

find the line that passes through the point (1,2) and is parallel to the line x+2y=3
-------------
Step 1, find the slope, m, of x+2y=3
------------
Parallel lines have the same slope.
Step 2, use
y-y1 = m*(x-x1) where (x1,y1) is the point (1,2)